An asset was purchased for $200,000 originally. Each year for the past 8 year, depreciation has been recorded which results in the asset having a current book value of $120,000. Now your firm is selling the asset for $140,000. What will be the total tax burden (or tax credit) caused by this sale? If it is a tax burden, enter the number as a positive number. If it is a tax credit, enter the number as a negative number. Assume a tax rate of 40%.
Answer #1

Book Value = 120,000

Selling Price = 140,000

Profit on sale = 140,000 - 120,000 = 20,000

Taxer = Profit on sales * Tax Rate

= 20,000 * 40%

= 8,000

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.